calculate the % sodium in a 0.721 g detergent sample


1
Expert's answer
2021-02-09T06:08:07-0500

Moles of NaP2O4 = 0.320 / 149 = 0.00215 mol

So, moles of Sodium is also = 0.00215 mol in NaP2O4

As we know according to law of conservation of mass,

Moles of sodium in detergent is also = 0.00215 mol

So, mass of Sodium = mole × molar mass = 0.00215 × 23 = 0.0494 g

%Sodium = (msodium × 100) / (mtotal) = 0.0494 × 100 / 0.721 = 6.85 %

% Sodium = 6.85 %